y= 2 +2 cos ( 3x+2pi) kindly solve this for me? thanks :-)
Answer by lwsshak3(11628)

y= 2 +2 cos ( 3x+2pi)
This is an equation for a cos function:
Its standard form: y=Acos(Bx-C), A=amplitude, Period=2π/B, Phase shift=C/B
For given function:y= 2 +2 cos ( 3x+2pi)
Amplitude=2
B=3
Period=2π/B=2π/3
C=2π
Phase shift: C/B=2π/3 (shift to the left)
Entire curve bumped up 2 units